CONSENT TO COPYRIGHT AGREEMENT

The Mathematical Journal of Okayama University hereby requests the Author(s) complete and return this form promptly so that the article may be readied for publication.

Title of Article:

Author(s):

The Mathematical Journal of Okayama University (“Journal”) and the above-named Author(s) agree that, effective upon acceptance for publication, the copyright of the above Article is transferred to the Journal throughout the world during the full term of copyright, including renewals and extensions, under the terms and conditions set forth below.

1. The Author(s) hereby consents that the Journal publishes the Article.

2. The Author(s) warrants that the Article has not been published before in any form except as a preprint, and that the Article is not concurrently submitted to another journal. The Author(s) agrees that the Journal shall have the right to publish the Article wholly or in part by any publication media, including but not limited to, electronic media and printed materials.

3. In publishing the Article, the Journal may alter the Article in order to unify the format of the Journal. The Author(s) shall be given an opportunity of making corrections, provided that the Author shall complete and return to the Journal the corrections by due date. The Journal may make further corrections on the proof of the Author(s) if necessary.

4. The Author(s) retains the right to keep the Article, after it has been published, in any preprint server approved by the Journal and/or in the Author’s own Web site in its published form (“Published Article”) for research, education and other academic purposes, provided that the Author(s) credits that it was published in the Journal, and that the copyright notice in proper form is placed on all copies. Any production or other form of reproduction not meeting these requirements shall be deemed to be unauthorized.

5. Upon request of the Author(s), the Journal may grant the Author(s), free of charge, the right to publish the Published Article or its revision for the purpose of publishing the complete works of the Author(s) and other publications of academic value. In each case of authorized reproduction, the Author(s) shall satisfy all procedural requirements of the editorial board of the Journal and ensure that the original publication by the Journal is properly credited.

6. Upon request of a third party in connection with an intended publication of academic value, the Journal may approve of a reproduction of the Published Article. Other matters pertaining to the Author’s copyright in the Article shall be governed by the Universal Copyright Convention as revised at Paris and the Copyright Law of Japan.

To be signed by one of the authors who agrees to inform the others, also on behalf of any co-authors:

Signature:

Date:

Name:

Address: